OpenStudy (philcua):

Will give medal Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (philcua):

can you help

OpenStudy (philcua):

@abb0t

OpenStudy (anonymous):

i think one of thems D

OpenStudy (anonymous):

i would say Band C

OpenStudy (anonymous):

Opps B and C

OpenStudy (philcua):

SO IT WOULD BE B AND C

OpenStudy (anonymous):

Ya i think so

geerky42 (geerky42):

B,C, and D. 0+0=0 So {0} is closed under addition.

OpenStudy (philcua):

So be B C AND D

geerky42 (geerky42):

Yep

OpenStudy (philcua):

oK THANK YOU CAN YOU HELP ME WITH SOME MORE

geerky42 (geerky42):

sure

OpenStudy (anonymous):

lol i got d right hahaha

OpenStudy (philcua):

Numbers p and q are positive integers. Which statements are always true? Choose all answers that are correct. A. p + q is a positive integer B. p – q is a negative integer C. p • q is a positive integer D. is a positive integer

OpenStudy (philcua):

If you refresh it will work

geerky42 (geerky42):

What do you think?

OpenStudy (philcua):

A and c

geerky42 (geerky42):

Right.

geerky42 (geerky42):

D is p/q ? I only see "is a positive integer"

OpenStudy (philcua):

Here the next one wait it is p/p is a positive integer that stands for fraction

OpenStudy (philcua):

@geerky42

OpenStudy (philcua):

That one is d

geerky42 (geerky42):

p/p ?

geerky42 (geerky42):

Then D is answer, because p/p is always 1, hence always positive integer.

OpenStudy (philcua):

yeah but like a fraction

OpenStudy (philcua):

Om thanks here is the last one

OpenStudy (philcua):

Is the set of multiples of 5 closed under division? Explain why or provide a counterexample if not. A. Yes, because the quotient of any two multiples of 5 is also a multiple of 5. B. Yes, because the quotient of any two multiples of 5 is also a multiple of 10. C. No, and a counterexample is 100 ÷ 5 = 20. D. No, and a counterexample is 25 ÷ 50 = .

OpenStudy (philcua):

@geerky42

geerky42 (geerky42):

What do you think?

OpenStudy (philcua):

on de there is a fraction of 1/2 as the answer

geerky42 (geerky42):

Yeah, D

OpenStudy (philcua):

I think it is b

OpenStudy (philcua):

But it is d

OpenStudy (philcua):

Are you there

geerky42 (geerky42):

Set of multiples of 5 = \(\{\ldots, -10, -5, 0, 5, 10,15,\ldots\}\) \(\dfrac{25}{50} = \dfrac{1}{2}\), right? Is 1/2 in set of multiples of 5?

OpenStudy (philcua):

yes it is 25 divided by 50

geerky42 (geerky42):

It is equal to 1/2 Is 1/2 divisible by 5?

OpenStudy (philcua):

Uh I dont think so

geerky42 (geerky42):

No, so since we can select two elements from set and produce new number that is not in set, this set is not closed under division, so answer is D.

geerky42 (geerky42):

Does that make sense?

OpenStudy (philcua):

So what do you think it is

OpenStudy (philcua):

And yes it does

geerky42 (geerky42):

Okay

OpenStudy (philcua):

Thank you so much